# Spectral Alignment in Forward–Backward Representations via Temporal Abstraction

### 20 Mar 2026 v1

@@ 1 Introduction
Figure 1: Q-function via Successor Representation (SR). The SR enables rapid value inference for arbitrary goals (star marker). Low-rank structure is desirable for navigation, as it preserves topological features (e.g., rooms) while suppressing transient dynamics. Top: In discrete MDPs, the SR can be computed from the transition matrix. Bottom: In continuous domains, Forward–Backward (FB) learning approximates the SR, where the embedding dimension controls the rank of the approximation. Low-rank structure can arise through (1) explicit constraints (e.g., SVD or small embeddings), (2) long horizons (high $\gamma$), or (3) temporal abstraction (e.g., action repetition). We show that, in continuous settings, temporal abstraction provides the spectral alignment needed for effective bootstrapping, whereas high $\gamma$ or overly restrictive bottlenecks can impair representation learning.
Scaling to continuous state–action spaces requires representations that are both expressive and computationally tractable. Classical formulations scale poorly with state-space size, motivating structured operator approximations. Forward–Backward (FB) representations address this challenge by learning a low-rank factorization of the SR directly from interaction (blier2021learning; touati2021learningone). By constraining the spectrum of the learned operator, FB aims to capture dominant long-horizon structure while discarding short-horizon dynamics. Yet a fundamental incompatibility exists: although FB enforces a low-rank constraint for tractability, the true SR in continuous environments often exhibits slow spectral decay. In such settings, high-frequency modes of the transition dynamics contribute significantly to the operator’s complexity (dubail2025shift).
@@ 2 Related Works
Successor representations were originally introduced as a task-agnostic predictive representation enabling rapid adaptation to new reward functions (dayan1993improvinggeneralization). More recent work has leveraged SR for transfer and zero-shot reinforcement learning (barreto2017SRfortransfer). However, computing the exact SR scales poorly with state-space dimensionality, motivating low-rank and parametric approximations that capture dominant long-horizon dynamics.
Forward-Backward representation learning methods (blier2021learning; touati2021learningone; touati2022doeszeroshot) address this challenge by learning factorizations of the SR that emphasize shared future state occupancies rather than fine-grained state distinctions. These methods implicitly assume that the SR admits a low-rank structure, but provide limited theoretical guidance on when such a structure should emerge. Our work complements FB learning by analyzing how the spectral properties of the transition dynamics, induced by the policy and environment, govern the effective rank of the SR and by proposing practical mechanisms that promote such structure.

### 23 Mar 2026 v2

@@ 1 Introduction
Figure 1: Q-function via Successor Representation (SR). The SR enables rapid value inference for arbitrary goals (star marker). Low-rank structure is desirable for navigation, as it preserves topological features (e.g., rooms) while suppressing transient dynamics. Top: In discrete MDPs, the SR can be computed from the transition matrix. Bottom: In continuous domains, forward-backward (FB) learning approximates the SR, where the embedding dimension controls the rank of the approximation. Low-rank structure can arise through (1) explicit constraints (e.g., SVD or small embeddings), (2) long horizons (high $\gamma$), or (3) temporal abstraction (e.g., action repetition). We show that, in continuous settings, temporal abstraction provides the spectral alignment needed for effective bootstrapping, whereas high $\gamma$ or overly restrictive bottlenecks can impair representation learning.
[…]
@@ 2 Related Works
[…]
SR were originally introduced as a task-agnostic predictive representation enabling rapid adaptation to new reward functions (dayan1993improvinggeneralization). More recent work has leveraged SR for transfer and zero-shot reinforcement learning (barreto2017SRfortransfer). However, computing the exact SR scales poorly with state-space dimensionality, motivating low-rank and parametric approximations that capture dominant long-horizon dynamics.
[…]
FB representation learning methods (blier2021learning; touati2021learningone; touati2022doeszeroshot) address this challenge by learning factorizations of the SR that emphasize shared future state occupancies rather than fine-grained state distinctions. These methods implicitly assume that the SR admits a low-rank structure, but provide limited theoretical guidance on when such a structure should emerge. Our work complements FB learning by analyzing how the spectral properties of the transition dynamics, induced by the policy and environment, govern the effective rank of the SR and by proposing practical mechanisms that promote such structure.
@@ Appendix A Appendix
Acknowledgments
This work was funded by the Carl Zeiss Foundation through the ReScaLe project.